# Remarks on the Stanley depth and Hilbert depth of monomial ideals with linear quotients

### v4

@@ 2. Main results
$\operatorname{depth}(S/I)=n-s-1,\text{ where }n-s=\min\{|Z_{j}|\;|\;1\leq j\leq m\}.$
Conjecture 2.8 is equivalent to the fact that there exists $i\in[n]$ such that there is no $1\leq j\leq m$ with $x_{i}\nmid u_{j}$, $x_{i}\in Z_{j}$ and $|Z_{j}|=n-s$.
Lemma 2.10.
Let $I\subset S$ be a monomial with linear quotients and $x_{i}$ a variable. Then $(x_{i},I)$ has linear quotients. Moreover, if $S^{\prime}=K[x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]$, then $(x_{i},I)=(x_{i},J)$, where $J\subset S^{\prime}$ is a monomial ideal with linear quotients.
Proof.
$((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})=(x_{i},\bar{Z}_{j_{k}}).$ (2.9)
Indeed, since $((u_{1},\ldots,u_{j_{k}-1}):u_{j_{k}})=(\bar{Z}_{j_{k}})$ and $x_{i}u_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})$ it follows that $(x_{i},\bar{Z}_{j_{k}})\subset((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})$. Conversely, assume that $v\in S$ is a monomial with $vu_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})=(x_{i},u_{1},\ldots,u_{j_{k}-1})$. If $x_{i}\nmid v$, then $vu_{j_{k}}\in(u_{1},\ldots,u_{j_{k}-1})$, hence $v\in(\bar{Z}_{j_{k}})$. If $x_{i}\mid v$, then $v\in(x_{i},u_{1},\ldots,u_{j_{k}-1})$. Hence the claim (2.9) is true and therefore $(x_{i},I)$ has linear quotients. Now, let $J=(u_{j_{1}},\ldots,u_{j_{p}})$. For any $k\geq 2$, we have that
@@
If Conjecture 2.11 is true and $I\subset S$ is a monomial ideal with linear quotients, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$.
@@
We use induction on $n\geq 1$. If $n=1$ then there is nothing to prove. Assume $n\geq 2$. Let $I\subset S$ be a monomial ideal with linear quotients and let $i\in[n]$ such that $\operatorname{depth}(I,x_{i})\geq\operatorname{depth}(I)$ and $\operatorname{sdepth}(S/(x_{i},I))\leq\operatorname{sdepth}(S/I)$. We consider the short exact sequence
@@
$\operatorname{sdepth}(S/I)\geq\operatorname{sdepth}(S/(I,x_{i}))=\operatorname{sdepth}(S^{\prime}/J)\geq$
@@ 3. Remarks on the Hilbert depth
Let $I=(u_{1},\ldots,u_{m})\subset S$ be a squarefree monomial with linear quotients, where $(u_{1},\ldots,u_{i-1}):u_{i}$ is generated by variables for any $2\leq i\leq m$. As we seen in the previous section, $I$ has a decomposition
@@
Moreoever, since $I$ is squarefree, $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{i}$ consists in the variables which are not in $(u_{1},\ldots,u_{i-1}):u_{i}$ it follows that $\operatorname{supp}(u_{i})\subset Z_{i}$ for all $1\leq i\leq m$. Therefore, if we denote $d_{i}=\deg(u_{i})$ and $n_{i}=|Z_{i}|$, then $d_{i}\leq n_{i}$, for all $1\leq i\leq n$.
@@
    $\alpha_{j}(I)=\sum\limits_{i=1}^{m}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq j\leq n$.
@@
    $\alpha_{j}(S/I)=\binom{n}{j}-\sum\limits_{i=1}^{m}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq j\leq n$.
@@
(1) For convenience, we assume that $u_{1}=x_{1}x_{2}\cdots x_{p}$ for some $p\leq n$. For $j\geq p$, a squarefree monomial of degree $j$ in $u_{1}K[Z_{1}]=u_{1}K[x_{1},\ldots,x_{n}]$ is of the form $v=uw$, where $w\in K[x_{p+1},\cdots,x_{n}]$ is squarefree of degree $j-p$. Hence, there are $\binom{n-j}{p-j}=\binom{n_{1}-j}{d_{1}-j}$ such monomials. Similarly, there are $\binom{n_{i}-j}{d_{i}-j}$ squarefree monomials of degree $j$ in $u_{i}K[Z_{i}]$ for all $2\leq i\leq m$. Hence, we get the required conclusion from (3.1).
@@
We recall the following combinatorial identity, see for instance [6, Lemma 2.5]:
@@
    $\beta_{k}^{d}(I)=\sum\limits_{i=1}^{r}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq k\leq d\leq n$.
@@
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}-\sum\limits_{i=1}^{r}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq k\leq d\leq n$.
By straightforward computations, we can deduce the identity
$\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{N-j}{D-j}=(-1)^{k}\binom{d}{k}\binom{N}{D}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-D}\\ {-d,-N}\end{matrix};{\displaystyle-1}\right],$ (3.3)
where ${}_{r}F_{s}\!\left[\begin{matrix}a_{1},\dots,a_{r}\\ b_{1},\dots,b_{s}\end{matrix};z\right]=\sum\limits_{k=0}^{\infty}\frac{(a_{1})_{k}\cdots(a_{r})_{k}}{k!\,(b_{1})_{k}\cdots(b_{s})_{k}}z^{k}$ is the standard hypergeometric notation; see [13] for futher details.
@@
    $\beta_{k}^{d}(I)=(-1)^{k}\binom{d}{k}\sum\limits_{i=1}^{r}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-d_{i}}\\ {-d,-n_{i}}\end{matrix};{\displaystyle-1}\right]$ for all $0\leq k\leq d\leq n$.
@@
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}+(-1)^{k+1}\binom{d}{k}\sum\limits_{i=1}^{r}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-d_{i}}\\ {-d,-n_{i}}\end{matrix};{\displaystyle-1}\right]$ for all $0\leq k\leq d\leq n$.
@@
Let $I\subset S$ be a squarefree monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$. Then
@@
Suppose that $\alpha_{n-1}(S/I)=s>0$. From [15, Lemma 2.1] we can assume that $\deg(u_{1})\geq\deg(u_{2})\geq\cdots\geq\deg(u_{m})$, where $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$ is the linear order on $G(I)$. If $u_{1}=x_{1}x_{2}\cdots x_{n}\in G(I)$ then $I=(u_{1})$ and $\operatorname{depth}(S/I)=n-1$, a contradiction. Therefore:
$\deg(u_{1})=\deg(u_{2})=\cdots=\deg(u_{s})=n-1\text{ and }\deg(u_{i})<n-1\text{ for }s+1\leq i\leq m.$
Let $L=x_{1}x_{2}\cdots x_{n}$. By reordering the variables, we can assume that
$u_{1}=L/x_{n},\;u_{2}=L/x_{1},\;u_{3}=L/x_{2},\ldots,u_{s}=L/x_{s-1}.$
By straightforward computations, it follows that:
$Z_{1}=\{x_{1},\ldots,x_{n}\},\;Z_{2}=Z_{1}\setminus\{x_{1}\},\ldots,\;Z_{s}=Z_{1}\setminus\{x_{s-1}\}.$
We claim that $s=m$. Indeed, if $s<m$ and $\deg(u_{s+1})<n-1$, then the ideal $(u_{1},\ldots,u_{s}):u_{s+1}$ is not generated by variables, a contradiction.
If $s=m=1$ then $I=(u_{1})$ and again we get a contradiction from the fact that $\operatorname{depth}(S/I)=n-1$ in this case. Hence $s\geq 2$ and
$I=(L/x_{n},L/x_{1},\ldots,L/x_{s-1}),\text{ where }L=x_{1}x_{2}\cdots x_{n}.$ (3.4)
From (3.4) and Lemma 3.3 it follows that
$\alpha_{j}(S/I)=\begin{cases}\binom{n}{j},&0\leq j\leq n-2\\ n-s,&j=n-1\\ 0,&j=n\end{cases}.$ (3.5)
From (3.5) we therefore get
$\beta_{n-1}^{n-1}(S/I)=\sum_{j=0}^{n-1}(-1)^{n-1-j}\alpha_{j}(S/I)\leq\sum_{j=0}^{n-1}(-1)^{n-1-j}\binom{n}{j}-2=1-2=-1,$
and therefore $\operatorname{hdepth}(S/I)\leq n-2$, as required. ∎
Remark 3.5.
Note that the inequality $\operatorname{hdepth}(S/I)\geq n-2$ follows also from (3.5), by straightforward computations, so it is not necessarily to use in the proof Theorem 2.4 and the fact that $\operatorname{hdepth}(S/I)\geq\operatorname{sdepth}(S/I)$. Also, from the proof of Theorem 3.4, if $I$ has at least one generator of degree $n-1$ then $I$ is generated in degree $n-1$.
We mention that the squarefree monomial ideals $I$ with linear quotients of degree $n-2$ were extensively studied in [10]. For such an ideal, using a similar argument as in the proof of Theorem 3.4, one can deduce that either $I$ is principal, either $\operatorname{hdepth}(S/I)=n-3$.

### v5

@@ 2. Main results
$\operatorname{depth}(S/I)=n-s-1,\text{ where }n-s=\min\{|Z_{j}|\;:\;1\leq j\leq m\}.$
We claim that Conjecture 2.8 is equivalent to the fact that there exists $i\in[n]$ such that there is no $1\leq j\leq m$ with $x_{i}\nmid u_{j}$, $x_{i}\in Z_{j}$ and $|Z_{j}|=n-s$. Indeed, with the notations of Lemma 2.9, if there is some $u_{j_{k}}$ with $x_{i}\nmid u_{j_{k}}$ and $x_{i}\in Z_{j_{k}}$ then $u_{j_{k}}K[Z_{j_{k}}\setminus\{x_{i}\}]$ is a subspace in the decomposition of the ideal with linear quotients $J\subset S^{\prime}=K[x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]$ and thus
$\operatorname{depth}(S/(I,x_{i}))=\operatorname{depth}(S^{\prime}/J)\leq(n-1)-s-1=n-s-2<\operatorname{depth}(S/I).$
The converse is similar.
@@
If Conjecture 2.11 is true and $I\subset S$ is a proper monomial ideal with linear quotients, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$.
@@
We use induction on $n\geq 1$. If $n=1$ then there is nothing to prove. Assume $n\geq 2$. Let $I\subset S$ be a monomial ideal with linear quotients and let $i\in[n]$ such that $\operatorname{depth}(S/(I,x_{i}))\geq\operatorname{depth}(S/I)$ and $\operatorname{sdepth}(S/(x_{i},I))\leq\operatorname{sdepth}(S/I)$. We consider the short exact sequence
@@
$\operatorname{sdepth}(S/I)\geq\operatorname{sdepth}(S/(I,x_{i}))=\operatorname{sdepth}(S^{\prime}/J)$
@@ 3. Remarks on the Hilbert depth
Let $I=(u_{1},\ldots,u_{m})\subset S$ be a proper squarefree monomial with linear quotients, where $(u_{1},\ldots,u_{i-1}):u_{i}$ is generated by variables for any $2\leq i\leq m$. As we seen in the previous section, $I$ has a decomposition
@@
Moreover, since $I$ is squarefree, $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and, for $2\leq i\leq m$, $Z_{i}$ consists in the variables which are not in $(u_{1},\ldots,u_{i-1}):u_{i}$, it follows that $\operatorname{supp}(u_{i})\subset Z_{i}$ for all $1\leq i\leq m$. Therefore, if we denote $d_{i}=\deg(u_{i})$ and $n_{i}=|Z_{i}|$, then $d_{i}\leq n_{i}$, for all $1\leq i\leq m$.
We use the convention $\binom{r}{s}=0$ for $s<0$.
@@
    $\alpha_{j}(I)=\sum\limits_{i=1}^{m}\binom{n_{i}-d_{i}}{j-d_{i}}$ for all $0\leq j\leq n$.
@@
    $\alpha_{j}(S/I)=\binom{n}{j}-\sum\limits_{i=1}^{m}\binom{n_{i}-d_{i}}{j-d_{i}}$ for all $0\leq j\leq n$.
@@
(1) For convenience, we assume that $u_{1}=x_{1}x_{2}\cdots x_{p}$ for some $p\leq n$. For $j\geq p$, a squarefree monomial of degree $j$ in $u_{1}K[Z_{1}]=u_{1}K[x_{1},\ldots,x_{n}]$ is of the form $v=u_{1}w$, where $w\in K[x_{p+1},\cdots,x_{n}]$ is squarefree of degree $j-p$. Hence, there are $\binom{n-p}{j-p}=\binom{n_{1}-d_{1}}{j-d_{1}}$ such monomials. Similarly, there are $\binom{n_{i}-d_{i}}{j-d_{i}}$ squarefree monomials of degree $j$ in $u_{i}K[Z_{i}]$ for all $2\leq i\leq m$. Hence, we get the required conclusion from (3.1).
@@
We recall the following combinatorial identity, which can be easily derived from the Chu-Vandermonde identity
@@
    $\beta_{k}^{d}(I)=\sum\limits_{i=1}^{m}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-d_{i}}{j-d_{i}}$ for all $0\leq k\leq d\leq n$.
@@
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}-\sum\limits_{i=1}^{m}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-d_{i}}{j-d_{i}}$ for all $0\leq k\leq d\leq n$.
If $k\geq D$ then, using (3.2) and taking $\ell=j-D$ we get
$\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{N-D}{j-D}=\sum_{\ell=0}^{k-D}(-1)^{k-D-\ell}\binom{d-D-\ell}{k-D-\ell}\binom{N-D}{\ell}=\binom{N-d+k-D-1}{k-D}.$ (3.3)
Note that (3.3) is trivially satisfied for $k<D$ also.
@@
    $\beta_{k}^{d}(I)=\sum\limits_{i=1}^{m}\binom{n_{i}-d+k-d_{i}-1}{k-d_{i}}$ for all $0\leq k\leq d\leq n$.
@@
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}-\sum\limits_{i=1}^{m}\binom{n_{i}-d+k-d_{i}-1}{k-d_{i}}$ for all $0\leq k\leq d\leq n$.
@@
Let $I\subset S$ be a proper squarefree monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$. Then
@@
Suppose that $\alpha_{n-1}(S/I)=s>0$. From [12, Lemma 2.1] we can assume that $\deg(u_{1})\leq\deg(u_{2})\leq\cdots\leq\deg(u_{m})$, where $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$ is the linear order on $G(I)$. If $m=1$ then $I=(u_{1})$ is principal, a contradiction with the hypothesis $\operatorname{depth}(S/I)=n-2$.
Note that, if $x_{1}x_{2}\cdots x_{n}\in G(I)$ then, since $I$ has linear quotients, it follows that $u_{1}=x_{1}x_{2}\cdots x_{n}$ and $I=(u_{1})$, a contradiction. Therefore
$\deg(u_{1})\leq\deg(u_{2})\leq\cdots\leq\deg(u_{m})\leq n-1.$
We claim that $\deg(u_{1})\geq s$. Assume by contradiction that $\deg(u_{1})=\ell<s$ and let’s say that $u_{1}=x_{1}\cdots x_{\ell}$. Then $v_{k}=x_{1}\cdots x_{n}/x_{k}\in I$ for all $\ell<k\leq n$ and thus $\alpha_{n-1}(S/I)\leq\ell$, a contradiction. In particular, we have $\alpha_{j}(S/I)=\binom{n}{j}$ for all $j\leq s-1$ and thus, from (1.1) and (3.2), it follows that
$\beta_{k}^{n-1}(S/I)=\sum_{j=0}^{k}(-1)^{k-j}\binom{n-1-j}{k-j}\binom{n}{j}=1\text{ for all }k\leq s-1.$ (3.4)
We assume by contradiction that $\operatorname{hdepth}(S/I)=n-1$. From (1.2), it follows that
$s=\alpha_{n-1}(S/I)=\sum_{j=0}^{n-1}\beta_{j}^{n-1}(S/I)\text{ with }\beta_{j}^{n-1}(S/I)\geq 0.$ (3.5)
Therefore, from (3.4) we get
$\beta_{j}^{n-1}(S/I)=0\text{ for all }s\leq j\leq n-1.$ (3.6)
From (1.2), (3.4) and (3.6) it follows that
$\alpha_{k}(S/I)=\sum_{j=0}^{k}\beta_{j}^{n-1}(S/I)\binom{n-1-j}{k-j}=\binom{n}{k}-\binom{n-s}{k-s}\text{ for all }0\leq k\leq n.$ (3.7)
From Lemma 3.1(2) and (3.7) it follows
$\sum\limits_{i=1}^{m}\binom{n_{i}-d_{i}}{s-d_{i}}=1.$ (3.8)
Since $s\leq d_{1}\leq d_{2}\leq\cdots\leq d_{m}$ and $d_{i}\leq n_{i}$ for all $1\leq i\leq m$, from (3.8) it follows that $d_{1}=s$ and $d_{i}>s$ for $2\leq i\leq m$. Since $n_{1}=n$, from Lemma 3.1(2) it follows that
$\alpha_{d_{2}}(S/I)=\binom{n}{d_{2}}-\sum\limits_{i=1}^{m}\binom{n_{i}-d_{i}}{d_{2}-d_{i}}\leq\binom{n}{d_{2}}-\binom{n-d_{1}}{d_{2}-d_{1}}-\binom{n_{2}-d_{2}}{0}=\binom{n}{d_{2}}-\binom{n-s}{d_{2}-s}-1,$
which contradicts (3.7). ∎
@@ Acknowledgements
We would like to express our gratitude to the anonymous referee who help us to correct, improve and clarify our manuscript.